- Baseplate and shield RTD temperature sensors are working, and digitally recorded. The temperature sensing diode on the PD mount seems to be unconnected.
- We made dark current vs bias measurements on the P6 and P2 elements at room temperature. We also checked P3 but it appears to be unconnected. Then, after realigning the P6 element, we measured its photocurrent vs bias at room temperature. The photocurrent and dark current data looks much like what Koji previously reported at room temperature.
- Going from 77 K to room temperature, the beam had to be moved downward to correct the alignment. The magnitude was about 1 clockwise turn of the vertical alignment knob on the steering mirror. The horizontal shift was very small.
- Armed with this knowledge, we re-wetted the cryostat with LN2, bringing the baseplate and shield temperatures down into the 80 K range. Our plan was to let it warm up overnight while running an automated test series of dark current and photocurrent vs bias and temperature. We offset the vertical alignment by a half turn, in hopes of having the PD well aligned near the expected QE sweet spot of 150-200 K.
- The procedure for adding LN2 to the cryostat involves temporarily disconnecting the fiber coupler. Sadly, it appears in our haste we did not fully recover the alignment after this step, or else the calibration somehow shifted by a lot. The photocurrent is at the 0.1 mA level instead of the ~1 mA we expect. Accordingly, this run will primarily be useful as a test of the automation, as an indication of dark current vs temperature, and as a very rough, qualitative indication of QE vs temperature.
- The script is running on qil-ws1 in a
screen session. It can be interrupted with
screen -RAad autorun to connect to the session, followed by Ctrl-C to kill the script.
- In the morning, we need to decide whether to open the cryostat and fix the wiring, or to repeat this run with tighter control over the alignment.